# Learning Online Belief Prediction for Efficient POMDP Planning in Autonomous Driving

### v1

@@ III-C Option-based MCTS Planner
where $s$ and $l$ represent the longitudinal and lateral directions along the reference path, $s(0)$, $l(0)$, $\dot{s}(0)$ are the initial state of the node, $\Delta\tau$ is the time interval, and $T_{o}$ is the option length. Then, we can transform the states in the Frenet frame back into the Cartesian frame $\left(x(\tau),y(\tau),v(\tau),a(\tau)\right)$ [30].
Objective Function. We calculate the cost of options considering collision risk, ride comfort, deviation from the route, and desired speed. The objective function of the planner is:
@@
$c_{r}(n)=\sum_{i}^{N}\sum_{j}^{M}\mathds{1}(p_{j}\geq p_{th})p_{j}\mathds{1}_{overlap}\left(\mathbf{s}^{e}(n),\mathbf{b}^{i}(j,n)\right),$ (7)
where $\mathbf{s}^{e}$ is the state of the ego agent determined by the generated options, $\mathbf{b}^{i}(j)$ is the $j$-th predicted trajectory of agent $i$, $p_{j}$ is the predicted likelihood of that trajectory. $\mathds{1}_{overlap}$ is a binary indicator to check whether the two trajectories collide, and $p_{th}$ is a threshold to filter out low probability futures.
@@
The prior $Pr(a)$ is given by the learned Q-value network $Q_{\epsilon}$, and we apply this prior to the root node, while setting the prior to a uniform distribution in other nodes. The sampling prior for the root node is calculated as:
@@
where $\mathbf{z}^{e}_{t}$ is the latent observation state for the ego agent at the current time, which is retrieved from the observation encoder and contains necessary information including its past states, other agents, and road maps.
@@ III-D Learning Framework
Online Learning. In this stage, we fix the weights of the observation encoder, and we learn the belief update model $f_{\rho}$ and fine-tune the trajectory decoder $f_{\phi}$ in a driving simulator. At the end of an episode, we obtain a prediction buffer and an observation buffer that contains the positions of all agents at each timestep. We utilize the GMM loss in Eq. 10 to minimize the prediction loss at each valid time step in hindsight. In addition, we leverage the clipped double Q-learning [31] to learn a Q-value network $Q_{\epsilon}$ to guide the tree search planner. The reward function from the environment for Q-learning is formulated as:
@@
We jointly learn the belief update and Q-network (with the MCTS planner) to improve the overall decision-making performance. The procedure for online belief update learning and Q-learning is summarized in Algorithm 2.
@@ IV-A Dataset and Simulator
We employ the Waymo Open Motion Dataset (WOMD) [32] to train the models and replay the traffic data for surrounding agents in the MetaDrive [33] simulator to conduct online training and testing. Additionally, we set the agent behavior to be reactive in the simulator, allowing other agents to perform basic reactive actions, such as avoiding collisions from the rear end of the ego vehicle. Fig. 3 shows some replayed scenarios from WOMD in the MetaDrive simulator and the obtained observation space for the ego agent.
Figure 3: Examples of replayed WOMD scenarios in the MetaDrive simulator and the processed observation space including surrounding map polylines and historical trajectories of agents.
@@ IV-B Implementation Details
MCTS Planner. The planning horizon is set to $T_{f}=80$ timesteps, with a time interval of $\Delta t=0.1\ s$, and the option length is set to $T_{o}=20$. This significantly reduces the search depth to $T_{d}=4$. We track the belief states of $N=8$ agents and set the probability threshold for calculating collision risk to $p_{th}=0.15$. The options at each node are derived from combinations of lateral velocity $v_{l}=[-1,0,1]\ m/s$ and longitudinal acceleration $a=[-4,-2,0,1,3]\ m/s^{2}$. Lateral velocity is only considered if the acceleration is $-2,0,1$, leading to a total of $11$ options. We also ensure that the velocity remains non-negative. The planner executes a total of $100$ search iterations and applies a discount factor of $\gamma=0.8$.
@@ IV-C Main Results
Baselines. In testing, we compare our method against three strong baseline approaches: RL (DQN) method [31], intelligent driver model (IDM) [34], and differentiable integrated prediction and planning (DIPP) [35]. In addition, we introduce several ablated versions of our approach. The offline model refers to the prediction model without online belief update, and it is only trained on the offline dataset. The online update model has two versions: one with the ego agent’s intentions in the online belief update, and one without. Furthermore, we evaluate the effectiveness of employing imitation learning (IL) (i.e., predicted future trajectory for the ego vehicle) [7, 8] and DQN as prior in the MCTS planner. In training, we introduce two baselines for comparison: DQN agent and MCTS planner with online belief update model but lacking Q-learning.
Metrics. The task-related metrics include success rate, average episodic reward, average episodic time (number of steps), and average log divergence (L2 distance) measuring the deviation of the ego vehicle from logged behavior. Regarding prediction performance, we employ three key metrics: minimum average displacement error (minADE), consistency (variation in predictions across consecutive frames), and score accuracy (whether the highest probability is correctly assigned to the closest modality to the ground truth). The prediction consistency is defined as:
@@
Method Success rate Task time Log divergence Episode reward Prediction minADE Consistency Score accuracy
@@
Testing Results. We evaluate our proposed method against other baselines in the testing scenarios, and the results are summarized in Table I. The key conclusions and findings are as follows. (1) The DQN agent shows the weakest performance, while both the rule-based IDM agent and the optimization-based DIPP method demonstrate favorable efficacy in testing. (2) The basic MCTS method without DQN guidance performs significantly worse, and DQN guidance is more effective than IL guidance, particularly when IL guidance is learned offline. (3) The prediction consistency of an offline-learned model is substantially worse compared to that of online belief update models. While offering similar prediction accuracy, these online belief update models significantly enhance prediction consistency and the accuracy of intention probability estimations, leading to improved decision-making performance. (4) Incorporating the ego vehicle’s intention into the belief update process in the online prediction model can further improve the decision-making performance in terms of the episode reward. Some qualitative results are presented in Fig. 5, which reveal that our proposed online update model provides more consistent and accurate trajectory predictions of interacting agents compared to the offline prediction model. The online prediction model is able to continuously improve its prediction accuracy (in terms of trajectory and probability) over time by receiving new observations. See closed-loop testing results.
@@ IV-D Ablation Study
Influence of option length. We investigate the impact of varying option lengths on the performance of the MCTS planner, as detailed in Table II. We find that shorter option lengths (1, 5, 10) lead to increased computation time and lower success rate and episode reward. This is because, within a limited number of search iterations, a deeper search depth makes it increasingly difficult to find the optimal action. Conversely, extending the option length to 40, which essentially simplifies the process into two-stage planning, yields a success rate comparable to an option length of 20, but the episode reward at this length is reduced. Consequently, an option length of 20 appears to provide a good balance between performance and computational efficiency.
@@ V Conclusions
We develop an online and closed-loop behavior prediction model for POMDP planning in autonomous driving. Specifically, we propose an online recurrent neural belief update model and an option-based MCTS planner guided by Deep Q-learning. The belief update model can incorporate the ego vehicle’s intention to capture the closed-loop interactions between the AV and other agents. We introduce an online learning framework, which combines belief update learning and deep Q-learning to guide tree search. We validate our framework in simulated environments based on real-world driving scenarios. The experimental results indicate that our proposed online belief update model can significantly improve the temporal consistency and accuracy of probability estimations. Furthermore, DQN guidance considerably boosts the efficacy of the MCTS planner, resulting in improved decision-making performance. Future research could extend this framework to broader POMDP settings, such as noisy observations, by learning to update the belief states of agents to account for both actual physical states and future intentions. Moreover, the option generation process can be improved by learning from expert data and by making the options more flexible in terms of their length and termination criteria.
Acknowledgement
This work was supported by Berkeley DeepDrive. We would like to extend our gratitude to Yixiao Wang for his insights and support, as well as Rouhollah Jafari and Stefano Bonasera from General Motors for their constructive feedback.

### v2

@@ III-C Macro-action-based MCTS Planner
where $s$ and $l$ represent the longitudinal and lateral directions along the reference path, $s(0)$, $l(0)$, $\dot{s}(0)$ are the initial state of the node, $\Delta\tau$ is the time interval, and $T_{o}$ is the macro-action length. Then, we can transform the states in the Frenet frame back into the Cartesian frame $\left(x(\tau),y(\tau),v(\tau),a(\tau)\right)$ [27].
Objective Function. We calculate the cost of macro-actions considering collision risk, ride comfort, deviation from the route, and desired speed. The objective function is:
@@
$\small c_{r}(n)=\sum_{i}^{N}\sum_{j}^{M}\mathds{1}(p_{j}\geq p_{th})p_{j}\mathds{1}_{overlap}\left(\mathbf{s}^{e}(n),\mathbf{b}^{i}(j,n)\right),$ (7)
where $\mathbf{s}^{e}$ is the state of the ego agent determined by the generated macro-actions, $\mathbf{b}^{i}(j)$ is the $j$-th predicted trajectory of agent $i$, $p_{j}$ is the predicted likelihood of that trajectory. $\mathds{1}_{overlap}$ is a binary indicator to check whether the two trajectories collide, and $p_{th}$ is a threshold to filter out low probability futures.
@@
DQN Guidance. DQN is used in MCTS as an action prior to bias exploring high-rewarding areas in the search tree. The prior $Pr(a)$ is given by the learned Q-value network $Q_{\epsilon}$, and we apply this prior to the root node, while setting the prior to a uniform distribution in other nodes. The sampling prior for the root node is calculated as:
@@
where $\mathbf{z}^{e}_{t}$ is the encoded observation state for the ego agent at the current time, which is retrieved from the observation encoder and contains necessary information including its past states, other agents, and road maps.
@@ III-D Learning Framework
Online Learning. In this stage, we fix the weights of the observation encoder and learn the belief update model $f_{\rho}$ and fine-tune the trajectory decoder $f_{\phi}$ in a driving simulator. At the end of an episode, we obtain a prediction buffer and an observation buffer that contains the positions of all agents at each timestep. We utilize the GMM loss in Eq. 10 to minimize the prediction loss at each valid time step in hindsight. In addition, we leverage the clipped double Q-learning [28] to learn a Q-value network $Q_{\epsilon}$ to guide the tree search planner. The reward function from the environment for Q-learning is formulated as:
@@
We jointly learn the belief update and Q-network (with the MCTS planner) to improve the overall decision-making performance. The procedure for online belief update model learning and Q-learning is summarized in Algorithm 2.
@@ IV-A Dataset and Simulator
We employ the Waymo Open Motion Dataset (WOMD) [29] to train the models and replay the traffic data for surrounding agents in the MetaDrive [30] simulator to conduct online training and testing. Additionally, we set the agent behavior to be reactive in the simulator, allowing other agents to perform basic reactive actions, such as avoiding collisions from the rear end of the ego vehicle. Fig. 3 shows a replayed scenario from WOMD in the MetaDrive simulator and the obtained observation space for the ego agent.
Figure 3: Example of replayed WOMD scenario in the MetaDrive simulator and the processed observation space including vectorized map polylines and historical trajectories of surrounding agents.
@@ IV-B Implementation Details
MCTS Planner. The planning horizon is set to $T_{f}=80$ timesteps, with a time interval of $\Delta t=0.1\ s$, and the option length is set to $T_{o}=20$. This significantly reduces the search depth to $T_{d}=4$. We track the belief states of $N=8$ agents and set the probability threshold for calculating collision risk to $p_{th}=0.15$. The options at each node are derived from combinations of lateral velocity $v_{l}=[-1,0,1]\ m/s$ and longitudinal acceleration $a=[-4,-2,0,1,3]\ m/s^{2}$. Lateral velocity is only considered if the acceleration is $-2,0,1$, leading to a total of $11$ options. We also ensure that the velocity remains non-negative. The planner executes a total of $100$ search iterations and applies a discount factor $\gamma=0.8$ and temperature parameter $C_{p}=100$ after careful tuning.
@@ IV-C Main Results
Baselines. In testing, we compare our method against several strong baseline approaches: RL (DQN) method [28], intelligent driver model (IDM) [31], and differentiable integrated prediction and planning (DIPP) [32]. We also integrate well-established trajectory prediction models from the WOMD benchmark, such as MTR [19] and GameFormer [33], into the MCTS planner to test their performance on planning. To ensure a fair comparison, we scale down these models to match the size of ours and train them using the same dataset.
Moreover, we introduce several ablated versions of our approach. The offline model denotes the prediction model that operates without online belief updates (i.e., decoding trajectories directly from encoded observations), and is only trained on the offline dataset. The proposed online update model has two variants: one incorporates the ego agent’s intentions in the online belief update, and the other does not. Furthermore, we evaluate the effectiveness of employing imitation learning (IL) (i.e., predicted future trajectory for the ego vehicle) [10, 11] and DQN as prior in the MCTS planner. For training comparisons, we introduce two baselines: a DQN agent and an MCTS planner that includes an online belief update model but omits Q-learning.
Metrics. The task-related metrics include success rate, average episode reward, task time (defined as the average number of steps per episode), and average log divergence (L2 distance) measuring the deviation of the ego vehicle from logged behavior. Regarding prediction performance, we employ three key metrics: minimum average displacement error (minADE), consistency (variation in predictions across consecutive frames), and score accuracy (whether the highest probability is correctly assigned to the closest modality to the ground truth). Prediction consistency is defined as:
@@
Method Success rate Task time Log divergence Episode reward Pred. minADE ($\downarrow$) Consistency ($\downarrow$) Score accuracy ($\uparrow$)
@@
Testing Results. We evaluate our proposed method against other baselines in the testing scenarios, and the results are summarized in Table I. The key conclusions and findings are as follows. (1) The DQN agent shows the weakest performance, while the rule-based IDM agent and the optimization-based DIPP method are more effective. Using established prediction models with our DQN-guided MCTS planner shows comparable performance to our offline-trained model but is less effective in planning than the online prediction model. (2) The basic MCTS method without DQN guidance performs significantly worse, and DQN guidance is more effective than IL guidance. (3) The prediction consistency of offline models is substantially worse than that of online belief update models. While offering similar prediction accuracy, the online belief update model significantly enhances prediction consistency and the accuracy of intention probability estimations, leading to improved decision-making performance. (4) Incorporating the ego vehicle’s intention into the belief update process in the online prediction model can further improve the decision-making performance in terms of the episode reward. Some qualitative results are presented in Fig. 5, which reveal that our proposed online update model provides more consistent and accurate trajectory predictions of interacting agents compared to the offline prediction model. The online prediction model can continuously improve its prediction accuracy (in terms of trajectory ADE and probability) by receiving new observations over time. The closed-loop testing results showcase the online prediction model’s much better prediction consistency.
@@ IV-D Ablation Study
Influence of macro-action length. We investigate the impact of varying macro-action lengths on the performance of the MCTS planner, as detailed in Table II. We find that shorter action lengths (1, 5, 10) lead to increased computation time and lower success rate and episode reward. This is because, within a limited number of search iterations, a deeper search depth makes it increasingly difficult to find better actions. Conversely, extending the macro-action length to 40, which essentially simplifies the process into two-stage planning, yields a success rate comparable to a macro-action length of 20, but the episode reward at this length is reduced. Consequently, a macro-action length of 20 appears to provide a good balance between performance and computational efficiency.
@@ V Conclusions
We develop an online behavior prediction model for POMDP planning in autonomous driving. Specifically, we propose a recurrent neural belief update model and a macro-action-based MCTS planner guided by deep Q-learning. We introduce an online learning framework, which combines belief update learning and deep Q-learning to guide tree search. We validate our framework in simulated environments based on real-world driving scenarios. The experimental results indicate that our proposed online belief update model can significantly improve temporal consistency and accuracy. Furthermore, DQN guidance considerably boosts the efficacy of the MCTS planner, resulting in improved decision-making performance. Future research could extend this framework to broader POMDP settings by learning to update both actual physical states and future intentions. Moreover, the macro-action generation process can be improved by learning from expert data and by making the macro-actions more flexible in terms of their length, manipulation, and termination criteria.
Acknowledgement. This work was supported by Berkeley DeepDrive. We would like to extend our gratitude to Yixiao Wang from UC Berkeley for his insights and support, as well as Rouhollah Jafari and Stefano Bonasera from General Motors for their constructive feedback.